[figure: 9.7 cm]

MAGNETOVOLUME EFFECT IN URANIUM COMPOUNDS

G.E. Grechnev^{a}, A.S. Panfilov^{a}, I.V. Svechkarev^{a}, O. Eriksson^{b}, B. Johansson^{b} and M. Mihalik^{c} aB. Verkin Institute for Low Temperature Physics, Kharkov, 310164, Ukraine

bCondensed Matter Theory Group, Department of Physics, University of Uppsala,

Box 530, S-751 21 Uppsala, Sweden cInstitute of Experimental Physics, 043 53, Košice, Slovakia

This contribution aims at finding an origin of pronounced volume effect on magnetic susceptibility χ in uranium compounds UCo_{2} and UX_{3} (X= Al, Si, Ga, Ge), situated well below and above the critical Hill limit, respectively. The volume dependent band structures of UX_{3} and UCo_{2} have been calculated in an external magnetic field by employing a relativistic full–potential LMTO method with the Zeeman term included for calculations of field–induced spin and orbital moments. The calculations have brought out strong p–f (for UX_{3}) and d–f (for UCo_{2}) hybridization, giving narrow bands of itinerant f states, as well as large and competing field–induced orbital and spin moments. The calculated volume derivatives of the induced moments appeared to be in agreement with the volume trends revealed for χ in UX_{3} compounds, as well as with dlnχ/dlnV derivatives evaluated from our experimental studies of the pressure effect on χ in UCo_{2} and UGa_{3}. The volume effect is found to be more pronounced for the induced orbital magnetic moments, due to the rapid quenching of these moments with increasing width of 5f bands under pressure. A large volume effect on the spin magnetic moment can not be entirely attributed to the volume dependence of the density of states at E_{F}.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : G.E. Grechnev

Address for correspondence : B. Verkin Institute for Low Temperature Physics, 47, Lenin’s pr., Kharkov, 310164, Ukraine

Email address : grechnev@ilt.kharkov.ua